why is -1*-1=+1?

Grade:11

1 Answers

vijay kumar pandey
13 years ago

well the question is not general.

if u r joking then it`s okay.otherwise i wanna make u more confused.

let x=1.

x^2=1,x^2-1=0

(x+1)(x-1)=0{since x=1,so x-1=0}

x+1*(x-1)=x-1

if u will cut x-1 both the sides,then u will get x+1=1.

now let`s put the inital assumed value of x that is 1,u know we get an cointradicting result,that is 1=2.

can u tell me the fault in this solution?